OpenStudy (anonymous):

proof: if three matrices \[A_{n \times 1}, B_{n \times n} , C_{n \times 1} \], then \[A^TBC=\sum_{i=1}^{n}B_{ij}[AC^T]_{ji}\]

OpenStudy (anonymous):

proof: if three matrices \[A_{n \times 1}, B_{n \times n} , C_{n \times 1} \], then \[A^TBC=\sum_{i=1}^{n}B_{ij}[AC^T]_{ji}\]

OpenStudy (anonymous):

@ganeshie8 @TuringTest @iambatman @Abhisar @iGreen

OpenStudy (mathmate):

The left hand side is a scalar, but the right hand side B and AC' are nxn. and the indices of AC' should probably be ij (instead of ji). Would the statement be: \(A^TBC=\sum_{i=1}^n\sum_{j=1}^{n}B_{ij}[AC^T]_{ij}\) or simply \(A^TBC=B_{ij}[AC^T]_{ij}\) using repeated indices

OpenStudy (michele_laino):

\[A ^{t}BC=a _{i}*b _{ij}*c _{j}=b _{ij}*(a _{i}*c _{j})=\] \[=b _{ij}*(A ^{t} C)_{ij}=b _{ij}*(C ^{t}A)_{ji}=B(C ^{t}A)\]

OpenStudy (michele_laino):

where I used the summation rule on repeated indices